if z1,z2,z3 are affixes of the vertices A,B,C respect. of a triangle ABC having centroid at G such that z=0 is the mid point of AG then....
(1)z1+z2+z3=0
(2)z1+4z2+z3=0
(3)z1+z2+4z3=0
(4)4z1+z2+z3=0

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear mann

Solution:

 

centroid of triangle G = (z1 + z2 +z3 )/3

given  [z1 + (z1 + z2 +z3 )/3]/2  =0

   so  4z1 + z2 +z3 =0

 

so 4th option is correct

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin